write the slope-intercept form of the equation of each line given the slope and y- intercept
1) slope=6, y-intercept=2
write the slope-intercept form of the equation of each line.
4) -3y=15-10x

Answer by richwmiller(17219)

y=mx+b
Where m is the slope and b is the y intercept.
1) y=6x+2
again the model is y=mx+b
4) -3y=15-10x
divide by -3
y=-5+10x
y=10x-5
where the slope is 10 and the y intercept is -5.
I removed the other examples because the site rules prohibit submitting more than one problem at a time .
plus no similar problems - you are supposed to learn from the problem we solve for you.
plus no more than 4 submissions daily